OpenStudy (anonymous):

3^(x+1) X 2^(x-2) =21

OpenStudy (alekos):

have you attempted this?

OpenStudy (isaiah.feynman):

I can't yet see through to the solution.

OpenStudy (alekos):

i have an answer x = log28/log6

OpenStudy (alekos):

anyone else?

OpenStudy (anonymous):

How do you do this?

OpenStudy (alekos):

is that the answer you have?

OpenStudy (alekos):

i'll have to go soon

OpenStudy (anonymous):

Nope,it's 1.86

OpenStudy (alekos):

1.86 happens to be log28/log6 on my calculator!!

OpenStudy (anonymous):

Alright,then how to do you do this question?

OpenStudy (alekos):

Its not easy but here it goes...... 1st step is to take the logs of both sides log[3^(x+1).2^(x-2)] = log21 then we get log[3^(x+1)] + log[2^(x-2)] = log21 follow so far?

OpenStudy (anonymous):

So we use product aw here?

OpenStudy (anonymous):

*law

OpenStudy (alekos):

yes that's right next step (x+1)log3 + (x-2)log2 = log21

OpenStudy (alekos):

now we expand and separate into like terms x(log3 + log2) = log21 + log4 - log3 happy with that?

OpenStudy (anonymous):

Yes,very much

OpenStudy (alekos):

ok now we get to the final result xlog6 = log(21x4/3) x = log28/log6 = 1.86 (2 sig figs)

OpenStudy (alekos):

there you go!

OpenStudy (anonymous):

Thanks a lot :D

OpenStudy (alekos):

no problem. you're welcome. i hope you really understand the steps

OpenStudy (anonymous):

I'll just take some time to let it all get in,I'll manage.